OpenStudy (anonymous):

The chance of winning on the type of lottery tickets we purchase is 0.05 or 1 in 20. We know to stop after we win, so this will be a geometric setting. We will play each Saturday until we win. What is the probability of winning on the second ticket?

OpenStudy (prettygirl_shynice):

This is a geometric distribution. X~Geo(0.05) So: 1 ticket P(x=1)=0.05 2 ticket: P(x=2)=0.95*0.05=0.0475 3 ticket: P(x=3)=0.95*0.95*0.05=0.045125 4 ticket: P(x=4)=0.95*0.95*0.95*0.05=0.04286875 and so on...

OpenStudy (anonymous):

Thanks :)

OpenStudy (prettygirl_shynice):

your welcome u got it

OpenStudy (anonymous):

It would be 1 in 400 or 0.0025 since it is 1/400 since it is 20 squares because you add 1 to the exponent every ticket. For example it would be 20 cubed third ticket

OpenStudy (anonymous):

Medal?